OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. x-intercept 8/13 and y-intercept 20/27

OpenStudy (lgbasallote):

i think i remember the equation \[\frac{x}{a} + \frac{y}{b} = 1\] where a is x-intercept and b is y-intercept...so substitute! \[\frac{x}{\frac{8}{13}} + \frac{y}{\frac{20}{27}} = 1\] simplify... \[\frac{13x}{8} + \frac{27y}{20} = 1\] now multiply everything by 40 to get rid of the denominators \[5(13x) + 2(27y) = 40\] \[65x + 54y = 40\] do you get it?

OpenStudy (anonymous):